OpenStudy (anonymous):

Vivian combined different amounts of white paint, blue paint, and green paint to make 145 milliliters (ml) of paint of a new shade. The amount of white paint was x ml. The amount of blue paint was 10 ml less than twice the amount of white paint, and the amount of green paint was half the amount of blue paint. This is represented in the equation below. x + (2x – 10) + one half (2x – 10) = 145 What was the difference in the amounts of white paint and blue paint Vivian combined? 35 ml 45 ml 30 ml 40 ml

OpenStudy (lgbasallote):

you should solve for x first

OpenStudy (anonymous):

yeah i did..but i kept getting a wrong answer

OpenStudy (lgbasallote):

x + 2x - 10 + 1/2(2x-10) = 145 x + 2x- 10 + x - 5 = 145 4x - 15 = 145 4x = 145 + 15 4x = 160 solve for x

OpenStudy (lgbasallote):

is that what you got?

OpenStudy (anonymous):

omg yes. but its wrong :/

OpenStudy (lgbasallote):

what did you get for x?

OpenStudy (anonymous):

70

OpenStudy (lgbasallote):

70? 4x = 160 divide both sides by 4 to get x

OpenStudy (anonymous):

or is it... x + (2x-10) + 1/2(2x-10)=145 x+(2x-10)+(1x-5)=145 x+(3x-15)=145 4x-15=145 4x=160 x=40

OpenStudy (lgbasallote):

x = 40 right

OpenStudy (anonymous):

Thank you for your help (:

OpenStudy (lgbasallote):

<tips hat>